OpenStudy (anonymous):

find the harmonic mean of 24 and 12

OpenStudy (lalaly):

\[\frac{2}{\frac{1}{24}+\frac{1}{12}} =\frac{2}{\frac{3}{24}} =\frac{48}{3}=16\]